OpenStudy (anonymous):

Use the Trapezoid below to help you answer the question. LM is the midsegment. HINT: LM=1/2(AB+DC) AB= 12 and LM= 5x-2. DC=8x-2. Find x.

OpenStudy (anonymous):

Here's the Trapezoid.

OpenStudy (anonymous):

sum up AB and DC then divide it by 2

OpenStudy (anonymous):

4x+5?

OpenStudy (anonymous):

Do we add everything to equal 360 or something like that?

OpenStudy (anonymous):

Or is that to find sides?

OpenStudy (anonymous):

Yeah kind of. I'm somewhat lost thought

OpenStudy (anonymous):

Help me to understand/solve 4x+5=12+8x−2/2

OpenStudy (anonymous):

If you dont mind

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

Okay so it will be, 8x+10/12+8x-2?

OpenStudy (anonymous):

both sides as in both 4x+5 and 12+8x-2?

OpenStudy (anonymous):

4x-5*2= 8x-10 12+8x-2*2= 24+16x-4

OpenStudy (anonymous):

oh okay I see

OpenStudy (anonymous):

And so then what do we do next

OpenStudy (anonymous):

Sorry I'm not sure how?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

okay lets start this again

OpenStudy (anonymous):